Me and John discussed a bit our next plans, then the facility lost
power.  Of  course I now assume we  are going to have beam back with 
snakes on somewhere on tuesday night. Last time this took about 14 hrs. 
                NO SHIFTS ARE BEING CANCELLED
_ 11/04 - C shift
   continue analysis of CC data, field ON/OFF. The mu-metal seems to 
   have worked pretty well with the TOF's the problem is that the CC
   rate is cut down a factor of 2. We have nevertherless install a 2nd
   layer of mumetal on the bottom tubes
_ 11/05 - A shift
  Continue internal CC gain matching with the flasher. The hall will be
  open but at some point it may need to be closed depending on the 
  progress in bringing the beam back. Of course beam has priority.
  Gain matching of CC adc-sum with cosmics.
_ 11/05 - B shift 
  Hall open. Finish installation of mu-metal (top TOF pmts). NC 
  trigger and gain matching with cosmic. Any CC issue leftover from A 
  shift
_ 11/05 - C shift
  Assuming beam comes on. optimize injection, find compton tune. Check 
  blast rates. Check tune with field on/off. Take Wch data with field on, 
  no lead sheets
_ 11/06 - A shift
  Assuming beam is on, eep data taking with field on.
_ 11/06 - B shift  Hall Open (?)
  any remaining cosmic/flasher issues. Connect flasher to NC and equalize 
  NC timing. Adjust NC delay to blat trigger (beam on). COME TO ANALYSIS 
  MEETING
If all this goes well, by thursday we may have TOFs, NC, CC calibrated and
in trigger, a known CC and TOF efficiency, a beam tune that allow parastic 
compton measurements. This is of course an ambitious plan, and of course is 
very dependable on beam recovery. 
ANALYSIS:
there is presently effort on a number of issues. I hope (and will push 
for) that by thursday we have a frozen commis/phase2 and **ALL** on-going
work goes to phase2/devel. All golden macros will be then backup in 
cvs and a description/user guide will be put in the blast book
Current issues are
  
  _ CC efficiency: everybody
  _ L/R asymmetry:  JS
  _ field on/off: (vz, tb)
  _ tdc_offsets: (pk)
  _ retime-OR (t0)
So that makes 5 fine new threads on Anaware. Note, we need tdc offsets 
also for the NC calibrations (cosmics). Of course we need the TOF first.
Other issues are the beamgate in hvdaemon (cc).
-- ________________________________________________________________________________ Tancredi Botto, phone: +1-617-253-9204 mobile: +1-978-490-4124 research scientist MIT/Bates, 21 Manning Av Middleton MA, 01949 ^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^
This archive was generated by hypermail 2.1.2 : Mon Feb 24 2014 - 14:07:28 EST